candygirl220:

write the solution in interval notation on 5/4(24-6m)>14-1/2(16-7m)

sillybilly123:

your answer?!?!

candygirl220:

yes

sillybilly123:

LOL, Yes is an acceptable answer?!?!

candygirl220:

what do mean?

sillybilly123:

i suspect there's a typo in yer answer .... or some other glitch .... and I would be surprised *and* delighted if the answer was a simple "Yes" good luck :)

sillybilly123:

photo the qu ?!?! and post

sillybilly123:

EG: \(\dfrac{5}{4(24-6m)} >14-\dfrac{1}{2(16-7m)}\) VS \(\dfrac{5}{4(24-6m)}>\dfrac{14-1}{2(16-7m)}\) VS \(\dfrac{5}{4(24-6m)} >14-\dfrac{1}{2} \cdot \dfrac{1}{16-7m}\) not clear which one it is Can think of other variations, too